A cylindrical drum has a diameter of 4 yards and a volume of 63 cubic cubic yards. What is the height of the drum? (Hint: Use V = пr2h)
(Points :5)

Answer by mananth(16949)

V= pi*r^2*h
r=2 yards
V= 63 cu.yards
..
63= pi*2^2*h
63/(4*pi)= h
height = 5.01 yards